# Energy Management for Renewable-Colocated Artificial Intelligence Data Centers

### v1

@@ Abstract
We develop an energy management system (EMS) for artificial intelligence (AI) data centers with colocated renewable generation. Under a profit-maximizing framework, the EMS of renewable-colocated data center (RCDC) co-optimizes AI workload scheduling, on-site renewable utilization, and electricity market participation. Within both wholesale and retail market participation models, the economic benefit of the RCDC operation is maximized. Empirical evaluations using real-world traces of electricity prices, data center power consumption, and renewable generation demonstrate significant profit gains from renewable and AI data center colocations.
Index Terms:
energy management system for AI data centers, large load colocaiton with distributed energy resources, flexible large load, AI data center scheduling.
@@ I Introduction
The escalating electricity consumption from data centers poses significant challenges for the power grid. Accommodating this surge requires substantial investments in grid expansion and generation capacity. However, such expansion typically involves long lead times, with some utilities reporting wait times of up to 7 to 10 years [2, 3]. From the perspective of the grid operation, colocating data centers with on-site renewable energy sources offers a promising approach to mitigating operational challenges. By enabling the direct use of locally generated renewable power, renewable colocation can reduce the net demand from data centers, alleviate network congestion, and advance decarbonization objectives.
How does renewable colocation benefit data centers? On-site generation of renewable reduces the cost of power consumption and carbon footprint, of course. However, will the cost reduction and environmental benefits offset the investment costs and the complexity of system integration for on-site renewable generation? To this end, it is essential to characterize the maximum benefits of renewable colocation, including not only cost savings from reduced electricity usage but also potential revenue from electricity market participation as a flexible demand resource, capable of providing ramping and regulation services, while taking into account that the AI data center has its priority serving AI tasks.
This paper focuses on the optimal energy management of a renewable-colocated data center (RCDC) by modeling the role of RCDC as a flexible prosumer participating in a wholesale or retail market while prioritizing the computational needs of its AI customers. The flexibility of AI data centers stems from the characteristics of two types of AI tasks: delay-tolerant (and thus deferrable) AI model-training tasks and delay-intolerant AI inferencing tasks that are non-deferrable; the former consumes 30-40% of annual energy consumption, according to the Electric Power Research Institute (EPRI) [4]. Leveraging the inherent flexibility of AI tasks, we develop a scheduling strategy that optimally allocates deferrable AI tasks to specific periods based on real-time renewable generation and electricity prices, thereby maximizing the economic benefits of RCDCs.
@@ I-B Contributions
The results presented here represent the first attempt to capture the full economic benefits of renewable-colocated data centers, featuring bidirectional participation in both wholesale and retail electricity markets. Our goal is to characterize the profit potential of data center-renewable colocation under a profit-maximizing framework that captures the power exchange dynamics between the RCDC and the grid, incorporating the operational characteristics of data center workloads.
@@
Second, we present a workload scheduling strategy that leverages the temporal characteristics of AI data center workloads. The EMS incorporates real-time on-site renewable generation and real-time electricity price information, allowing deferrable tasks to be strategically delayed to maximize profits from task processing and renewable energy sales.
@@
TABLE I: Decision Variables and System Parameters
@@ I-C Notations
The notations used in this paper follow standard conventions. We use $x$ to represent a scalar and $\bm{x}$ for a vector. We define $[x]:=\{1,\cdots,x\}$. For future reference, key variables and system parameters are summarized in Table I.
@@ II-A RCDC Operation Model
where $W_{t}$ denotes the amount of task processed at time $t$. $K$ denotes the number of segments in the piecewise linear function. Each segment is characterized by a pair of parameters $(\alpha_{k},\beta_{k})$, with $\{(\alpha_{k},\beta_{k})\}_{k\in[K]}$ specifying the slope and intercept. The corresponding segment domains are denoted by $\mathcal{P}_{k}^{\text{\tiny D}}$. The indictor function $\mathds{1}_{\{P_{t}^{\mbox{\tiny D}}\in\mathcal{P}_{k}^{\mbox{\tiny D}}\}}$ equals one if the power input falls within segment $k$, and zero otherwise.
The concave piecewise linear function effectively captures the characteristic of diminishing efficiency, i.e., as power input increases, the computational efficiency of the data center declines. Given that the data center comprises multiple servers, potentially exhibiting heterogeneous efficiencies, the piecewise linear approximation provides a suitable aggregate representation of the data center’s efficiency profile. In practice, (1) can be derived from empirical workload and power consumption data.
@@
Retail market pricing is regulated under a specific tariff for the RCDC as a commercial customer. A standard tariff applied by most distribution utilities is the net energy metering (NEM) tariff, which prices customer imports and exports differently. In particular, energy drawn from the grid costs the RCDC $\pi_{t}^{+}$ (\$/kWh) at $t$, and energy exported to the grid costs $-\pi_{t}^{-}$ (\$/kWh). To recover infrastructure-related costs, utilities commonly impose demand charges at $\lambda$ (\$/kW) for RCDCs as large loads. RCDC may also be subject to other fixed (connection) charges, which are ignored in our formulation since they do not affect RCDC scheduling decisions.
@@ II-B RCDC Workload Scheduling
III Profit-Maximizing Operation
The RCDC generates revenue through task processing and by exporting excess renewable power to the grid, while incurring costs from purchasing grid power when on-site generation is insufficient. We assume that the data center does not drop any non-deferrable tasks and that the deadline of deferrable tasks is strict [^2] For simplicity, we consider a single set of deferrable tasks. The model can be readily extended to accommodate multiple sets with varying deadlines.. Accordingly, the RCDC’s operational profit over the horizon $T$ is determined by the total revenue from surplus renewable exports minus the total cost of grid electricity purchases. Here we exclude the revenue from AI customers for processing their AI tasks, as such revenue are independent of EMS decisions provided that customer requests are satisfied.
Specifically, the objective function $J_{\mbox{\tiny T}}(\bm{P}^{\mbox{\tiny EX}},\bm{P}^{\mbox{\tiny IM}};\bm{\pi})$ of the EMS is defined by the profit from RCDC’s electricity market participation
$J_{\mbox{\tiny T}}(\bm{P}^{\mbox{\tiny EX}},\bm{P}^{\mbox{\tiny IM}};\bm{\pi})=J_{\mbox{\tiny T}}^{\mbox{\tiny EX}}(\bm{P}^{\mbox{\tiny EX}};\bm{\pi})-J_{\mbox{\tiny T}}^{\mbox{\tiny IM}}(\bm{P}^{\mbox{\tiny IM}};\bm{\pi}),$ (2)
where $\bm{P}^{\mbox{\tiny EX}}=\{P_{t}^{\mbox{\tiny EX}}\}_{t=1}^{T}$ represents the power exported to the grid, and $\bm{P}^{\mbox{\tiny IM}}=\{P_{t}^{\mbox{\tiny IM}}\}_{t=1}^{T}$ denotes the power imported from the grid. The vector $\bm{\pi}$ represents the pricing parameters of the wholesale or retail markets, as defined below.
@@ III Profit-Maximizing Operation
The profit-maximization problem is formulated as follows:
$\underset{\{\bm{P}^{\mbox{\tiny EX}},\bm{P}^{\mbox{\tiny IM}},\bm{P}^{\mbox{\tiny D}},\bm{W}^{\mbox{\tiny DF}}\}}{\rm maximize}$ $J_{\mbox{\tiny T}}(\bm{P}^{\mbox{\tiny EX}},\bm{P}^{\mbox{\tiny IM}};\bm{\pi})$ (3a)
@@
Electricity Cost (Percentage Savingsa ) \$$3.82\times 10^{6}$$1.19\times 10^{6}$ (68.8%) \$$0.78\times 10^{6}$ (79.5%)
Investment-Adjusted Savings \$$0.17\times 10^{6}$$0.58\times 10^{6}$
@@
Electricity Cost (Percentage Savingsa ) \$$4.88\times 10^{6}$$2.04\times 10^{6}$ (58.2%) \$$1.73\times 10^{6}$ (64.5%)
Investment-Adjusted Savings \$$0.38\times 10^{6}$$0.69\times 10^{6}$
@@
    Percentage of cost savings compared to the no colocation configuration.
@@ IV Numerical Study
We assumed that 40% of the workload was deferrable. Table II summarizes RCDC’s monthly operational outcomes under both wholesale and retail settings. The no colocation configuration incurred the highest costs, whereas the optimal colocation configuration achieved the lowest, due to the combined benefits of renewable utilization and workload scheduling [^4] The presented electricity cost is the net cost, which accounts for revenue from selling renewables to the grid.. In both markets, the optimal colocation configuration demonstrated higher market participation than the colocation configuration, as the volatility of electricity prices allowed for strategic buying grid electricity during low-price periods and selling renewables during high-price intervals. In the retail market, peak demand was effectively managed, resulting in a lower demand charge.
We calculated the amortized monthly cost of the renewable generation’s capital and operational expenditures, assuming a 30-year lifespan. This amounted to approximately \$2.46 million per month for 150 MW of land-based wind generation [19]. The cost savings from renewable colocation could offset this amount, supporting the economic viability of renewable-colocated data centers.
Figure 1: Electricity cost savings achieved by the optimal colocation configuration compared to the no colocation configuration under varying fractions of deferrable tasks.
Figure 2: Electricity cost savings achieved by the optimal colocation configuration compared to the colocation configuration under varying fractions of deferrable tasks.
Figs. 1-2 illustrate how the share of deferrable tasks influences cost savings achieved by RCDC’s optimal EMS. In the wholesale market, a higher proportion of deferrable tasks enabled the data center to shift more workload to periods with lower electricity prices, resulting in greater cost savings. However, the marginal benefit diminished at higher deferrable ratios due to constraints imposed by facility capacity and the availability of renewable generation. In the retail market, cost reductions were smaller and less sensitive to the deferrable task fraction. This is primarily because the need to avoid high demand charges imposed stricter constraints on the scheduling flexibility of deferrable tasks.
We further analyzed the effect of the renewable-to-data center capacity ratio, as shown in Figs. 3-4. As this ratio increased, the renewable supply became more adequate to meet demand, thereby enhancing the cost advantage of configurations with colocated renewables over the no colocation configuration. As illustrated in Fig. 4, the benefit of workload scheduling in the optimal colocation configuration remained constant in the wholesale market as renewable capacity increased. This is because, relative to the colocation configuration, the additional benefit stemmed solely from exploiting fluctuations in LMPs. In contrast, in the retail market, the EMS also considers reducing peak demand. Since different levels of renewable generation result in different electricity import patterns, the benefit from scheduling varies accordingly.
@@
Figure 3: Electricity cost savings achieved by the optimal colocation configuration compared to the no colocation configuration under varying renewable-to-data center capacity ratios.
@@
Figure 4: Electricity cost savings achieved by the optimal colocation configuration compared to the colocation configuration under varying renewable-to-data center capacity ratios.
@@ V Conclusion
To enhance the economic performance of the RCDC, we introduce an optimization framework that models the interaction between workload execution and power exchange with the grid. By strategically scheduling deferrable tasks based on real-time renewable generation and electricity prices, the RCDC can maximize profit. As electricity costs become a major component of data center expenses, this approach offers a promising solution for improving cost-efficiency.

### v2

@@ Abstract
We develop an energy management system (EMS) for artificial intelligence (AI) data centers with colocated renewable generation. Under a cost-minimizing framework, the EMS of renewable-colocated data center (RCDC) co-optimizes AI workload scheduling, on-site renewable utilization, and electricity market participation. Within both wholesale and retail market participation models, the economic benefit of the RCDC operation is maximized. Empirical evaluations using real-world traces of electricity prices, data center power consumption, and renewable generation demonstrate significant electricity cost reduction from renewable and AI data center colocations.
@@ I Introduction
The escalating electricity demand from data centers poses significant challenges for the power grid. Accommodating this surge requires substantial investments in grid expansion and generation capacity. However, such expansion typically involves long lead times, with some utilities reporting wait times of up to 7 to 10 years [2, 3]. From the perspective of the grid operation, colocating data centers with on-site renewable energy sources offers a promising approach to mitigating operational challenges. By enabling the direct use of locally generated renewable power, renewable colocation can reduce the net demand from data centers, alleviate network congestion, and advance decarbonization objectives.
How does renewable colocation benefit data centers? On-site generation of renewable reduces the cost of power consumption and carbon footprint, of course. However, will the cost reduction and environmental benefits offset the investment costs and the complexity of system integration for on-site renewable generation? To this end, it is essential to characterize the benefits of renewable colocation in terms of cost savings, arising from both reduced grid electricity usage and revenues from electricity market participation that offset operational costs, while ensuring that the AI data center prioritizes serving AI tasks.
This paper focuses on the optimal energy management of a renewable-colocated data center (RCDC) by modeling the role of RCDC as a flexible prosumer participating in a wholesale or retail market while prioritizing the computational needs of its AI customers. The flexibility of AI data centers stems from the characteristics of two types of AI tasks: delay-tolerant (and thus deferrable) AI model-training tasks and delay-intolerant (non-deferrable) AI inferencing tasks, with the former accounting for 30-40% of annual energy consumption, according to the Electric Power Research Institute (EPRI) [4]. Leveraging the inherent flexibility of AI tasks, we develop a scheduling strategy that optimally allocates deferrable AI tasks to specific periods based on real-time renewable generation and electricity prices, thereby maximizing the economic benefits of RCDCs.
@@ I-B Contributions
The results presented here represent the first attempt to capture the full economic benefits of renewable-colocated data centers, featuring bidirectional participation in both wholesale and retail electricity markets. Our goal is to characterize the cost-minimization potential of data center-renewable colocation under a framework that captures the power exchange dynamics between the RCDC and the grid, incorporating the operational characteristics of data center workloads.
@@
Second, we present a workload scheduling strategy that leverages the temporal characteristics of AI data center workloads. The EMS incorporates real-time on-site renewable generation and electricity price information, allowing deferrable tasks to be strategically delayed to reduce net costs through optimized task processing and renewable energy sales.
@@
TABLE I: Variables and System Parameters
@@ I-C Notations
The notations used in this paper follow standard conventions. We use $x$ to represent a scalar and $\bm{x}$ for a vector. We define $x]:=\{1,\cdots,x\}$. For future reference, key variables and system parameters are summarized in Table I.
@@ II-A RCDC Operation Model
where $W_{t}$ denotes the amount of task processed at time $t$. $K$ denotes the number of segments in the piecewise linear function. Each segment is characterized by a pair of parameters $(\alpha_{k},\beta_{k})$, with $\{(\alpha_{k},\beta_{k})\}_{k\in[K]}$ specifying the slope and intercept. The corresponding segment domains are denoted by $\mathcal{P}_{k}^{\text{\tiny D}}$. The indictor function $\mathds{1}_{\{P_{t}^{\mbox{\tiny D}}\in\mathcal{P}_{k}^{\mbox{\tiny D}}\}}$ equals one if the power input falls within segment $k$, and zero otherwise. The piecewise linear formulation, adopted in prior work such as [16], provides a tractable and effective representation of the data center’s efficiency profile. In practice, (1) can be derived from empirical workload and power consumption data.
@@
Retail market pricing is regulated under a specific tariff for the RCDC as a commercial customer. A standard tariff applied by most distribution utilities is the net energy metering (NEM) tariff, which prices customer imports and exports differently. In particular, energy drawn from the grid costs the RCDC $\pi_{t}^{+}$ (\$/kWh) at $t$, while energy exported to the grid yields a payment of $\pi_{t}^{-}$ (\$/kWh). To recover infrastructure-related costs, utilities commonly impose demand charges at $\lambda$ (\$/kW) for RCDCs as large loads. RCDC may also be subject to other fixed (connection) charges, which are ignored in our formulation since they do not affect RCDC scheduling decisions.
@@ II-B RCDC Workload Scheduling
III Cost-Minimizing Operation
The RCDC generates revenue through task processing and by exporting excess renewable power to the grid, while incurring costs from purchasing grid power when on-site generation is insufficient. We assume that the data center does not drop any non-deferrable tasks and that the deadline of deferrable tasks is strict [^2] For simplicity, we consider a single set of deferrable tasks. The model can be readily extended to accommodate multiple sets with varying deadlines.. Accordingly, for the purpose of EMS optimization, we exclude the revenue from AI customers, as it is independent of EMS decisions provided that customer requests are satisfied, and focus solely on the operational net cost associated with electricity market interactions.
Specifically, the EMS objective function $J_{\mbox{\tiny T}}(\bm{P}^{\mbox{\tiny IM}},\bm{P}^{\mbox{\tiny EX}};\bm{\pi})$ is defined as the RCDC’s net electricity cost, determined by electricity prices, grid purchases, and surplus renewable exports.
$J_{\mbox{\tiny T}}(\bm{P}^{\mbox{\tiny IM}},\bm{P}^{\mbox{\tiny EX}};\bm{\pi})=J_{\mbox{\tiny T}}^{\mbox{\tiny IM}}(\bm{P}^{\mbox{\tiny IM}};\bm{\pi})-J_{\mbox{\tiny T}}^{\mbox{\tiny EX}}(\bm{P}^{\mbox{\tiny EX}};\bm{\pi}),$ (2)
where $\bm{P}^{\mbox{\tiny IM}}=\{P_{t}^{\mbox{\tiny IM}}\}_{t=1}^{T}$ denotes the power imported from the grid, and $\bm{P}^{\mbox{\tiny EX}}=\{P_{t}^{\mbox{\tiny EX}}\}_{t=1}^{T}$ represents the power exported to the grid. The vector $\bm{\pi}$ represents the pricing parameters of the wholesale or retail markets, as defined below.
@@ III Cost-Minimizing Operation
The cost-minimization problem is formulated as follows:
$\underset{\{\bm{P}^{\mbox{\tiny EX}},\bm{P}^{\mbox{\tiny IM}},\bm{P}^{\mbox{\tiny D}},\bm{W}^{\mbox{\tiny DF}}\}}{\rm minimize}$ $J_{\mbox{\tiny T}}(\bm{P}^{\mbox{\tiny IM}},\bm{P}^{\mbox{\tiny EX}};\bm{\pi})$ (3a)
@@
Electricity Cost (Percentage Reductiona ) $3.82\times 10^{6}$ $1.19\times 10^{6}$ (68.8%) $0.78\times 10^{6}$ (79.5%)
Investment-Adjusted Cost Reduction $0.17\times 10^{6}$ $0.58\times 10^{6}$
@@
Electricity Cost (Percentage Reductiona ) $4.88\times 10^{6}$ $2.04\times 10^{6}$ (58.2%) $1.73\times 10^{6}$ (64.5%)
Investment-Adjusted Cost Reduction $0.38\times 10^{6}$ $0.69\times 10^{6}$
@@
    Percentage of cost reduction compared to the no colocation configuration.
@@ IV Numerical Study
IV-A Colocation Cost Benefits
Table II summarizes RCDC’s monthly operational outcomes under both wholesale and retail settings, assuming 40% deferrable workload. Compared to the no colocation configuration, we observed 68.8% cost reduction in the wholesale market and 58.2% in the retail market under the colocation configuration, while the optimal colocation configuration achieved reductions of 79.5% and 64.5%, respectively [^4] The presented electricity cost is the net cost, which accounts for revenue from selling renewables to the grid.. According to the wind energy cost report published by the National Renewable Energy Laboratory (NREL), the capital expenditure for utility-scale, land-based wind projects is \$1,968/kW, with operational expenditures of \$43/kW/year in 2023 [20]. Assuming a 30-year lifespan and a corresponding monthly mortgage interest rate of 0.564% [^5] This interest rate reflects typical US commercial bank lending rates [21]., the amortized monthly cost for a 150 MW land-based wind installation is approximately \$2.46 million. The cost reduction from renewable colocation could offset this amount, supporting the economic viability of renewable-colocated data centers.
The benefit of colocation arose from reduced grid electricity purchases, while optimal colocation further enhanced savings by scheduling deferrable tasks and strategically participating in electricity markets. Compared to the simple RCDC, the optimal configuration exhibited greater market participation, buying grid electricity during low-price periods and selling renewables at higher prices. In the retail market, it additionally mitigated peak demand, thereby lowering the demand charge.
IV-B Impact of Flexibility on Cost Reduction
Figs. 1-2 indicate that RCDC’s optimal EMS achieved larger cost reductions as the share of deferrable tasks increased, with diminishing marginal benefits at higher fractions. This trend was more pronounced in the wholesale market, whereas cost reductions in the retail market were less sensitive to the deferrable task fraction.
Figure 1: Investment-adjusted cost reduction in the wholesale market under varying fractions of deferrable tasks.
Figure 2: Investment-adjusted cost reduction in the retail market under varying fractions of deferrable tasks.
In both markets, a higher proportion of deferrable tasks enabled the data center to shift more workload to periods with lower electricity prices, enhancing savings under the optimal colocation configuration. By contrast, in the colocation configuration, cost reduction remained constant regardless of the deferrable task fraction, as workload scheduling was not exploited and economic gains arose solely from renewable energy utilization.
IV-C Impact of Renewable Generation on Cost Reduction
Figs. 3-4 demonstrate that renewable colocation consistently yielded positive investment-adjusted cost reductions across all renewable-to-data center capacity ratios, compared to the no colocation configuration. Moreover, the cost savings increased with the ratio, indicating that larger colocated renewable capacity strengthened the economic advantage. This effect resulted from higher renewable availability more effectively offset grid purchases. Although larger renewable deployments incurred greater capital investment, the resulting cost reductions proved sufficient to ensure net economic benefits.
@@ IV-C Impact of Renewable Generation on Cost Reduction
Figure 3: Investment-adjusted cost reduction in the wholesale market under varying renewable-to-data center capacity ratios.
@@
Figure 4: Investment-adjusted cost reduction in the retail market under varying renewable-to-data center capacity ratios.
@@ V Conclusion
We introduce an optimization framework that models the interaction between workload execution and grid power exchange. By strategically scheduling deferrable tasks based on real-time renewable generation and electricity prices, the RCDC can minimize its net electricity costs. As electricity costs become a major component of data center expenses, this approach offers a promising solution for improving cost-efficiency.